OpenStudy (grnsntcrtvclr):

How do I evaluate cos(1) without a calculator? I know cos(0), because zero degrees on the unit circle has the coordnate (1,0), but I don't see a 1 degree or radian on the unit circle anywhere??

OpenStudy (unklerhaukus):

evaluate in terms of what exactly?

OpenStudy (mhchen):

he means in terms of radians.

OpenStudy (dumbcow):

Use a Taylor series to approximate the cosine function \[\cos x = 1 - \frac{x^2}{2} + \frac{x^4}{24} -\frac{x^6}{720} ...\] Make sure you use radians, so if you want cos of 1 degree, plug in x = (pi/180) Of course, you have to use your calculator to evaluate the polynomial as well ...